Coronado Company purchased equipment for $216,240 on October 1, 2017. It is estimated that the equipment will have a useful life of 8 years and a salvage value of $12,240. Estimated production is 40,000 units and estimated working hours are 20,300. During 2017, Coronado uses the equipment for 530 hours and the equipment produces 1,100 units.

Compute depreciation expense under the following methods. Coronado is on a calendar-year basis ending December 31. (Round rate per hour and rate per unit to 2 decimal places, e.g. 5.35 and final answers to 0 decimal places, e.g. 45,892.)

Straight-line method for 2017

Sum-of-the-years'-digits method for 2019

Double-declining-balance method for 2018
Answer #1

1) Straight line method for 2017 = (216240-12240/8)*3/12 = 6375

2) Sum of year digit :

Sum of year digit = 8+7+6+5+4+3+2+1 = 36

Depreciation expense for 2019 = (204000*7/36*9/12+204000*6/36*3/12) = 38250

3) Double decline balance

Depreciation expense for 2018 = (216240*25%*9/12+216240*75%*25%*3/12) = 50681